OpenStudy (anonymous):

Please Explain: Solve for x in the nearest thousandth log(x+2) - log 2x= 8

Nnesha (nnesha):

log properties are you familiar with log properties ?

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

so 1st step) condense log at left side

OpenStudy (anonymous):

How would we do the first step? Are we isolating the x?

OpenStudy (anonymous):

(I'm not too familiar with logs, but a problem like this will be on my final)

Nnesha (nnesha):

okay first step would be condense log at left side look at the log properties which one you should apply ?

Nnesha (nnesha):

don't worry log are easy :-) you will be fine

Nnesha (nnesha):

key: there is negative sign in your question so you can change that to what ?(according to the log properties )?

OpenStudy (anonymous):

would it be the quotient rule?

Nnesha (nnesha):

yes right :-)

Nnesha (nnesha):